OpenStudy (anonymous):

Find the two unit vectors in R^2 that are perpendicular to the line y=-2

OpenStudy (anonymous):

Consider any two points on \(y\). Let them be \(P_1\) and \(P_2\). \(P_2-P_1\) is the directional vector of \(y\).

OpenStudy (anonymous):

So for example... \((0,-2)\) and \((1,-2)\).

OpenStudy (anonymous):

\[ (1,-2)-(0,-2) = \langle 1,0\rangle \]

OpenStudy (anonymous):

We can find perpendicular vectors based on the fact that the dot product will equal 0.

OpenStudy (anonymous):

so is <1,0> the vector that represents y=-2? or is that one of my perpendicular vectors

OpenStudy (anonymous):

It represents the direction of \(y=-2\).

OpenStudy (anonymous):

Consider the vector: \(\langle a,b\rangle\)

OpenStudy (anonymous):

\[ 0 = \langle a,b\rangle \cdot \langle 1,0\rangle = a+0=a \]

OpenStudy (anonymous):

Okay that makes sense. Sorry if I sound completely stupid with this. I'm fairly new to vectors. My professor basically showed us how to add and subtract vectors and dabbled into multiplying them then assigned us this stuff without explaining any of it. So I'm completely lost.

OpenStudy (anonymous):

So \(a=0\). Since it is a unit vector then: \[ 1^2 = \langle a,b\rangle \cdot \langle a,b\rangle = a^2+b^2 = 0^2+b^2 = b^2\implies b^2=1^2=1 \]

OpenStudy (anonymous):

Now, do you know the solutions to: \[ b^2 = 1 \]?

OpenStudy (anonymous):

No, you completely lost me

OpenStudy (anonymous):

Do you know how to compute: \[ \langle a,b\rangle \cdot \langle a,b\rangle \]?

OpenStudy (anonymous):

a^2+b^2 ?

OpenStudy (anonymous):

Yes. Now we already established that \(a=0\), in order to make the vector perpendicular.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Now, do you know what a unit vector is?

OpenStudy (anonymous):

a vector with the length of 1

OpenStudy (anonymous):

?

OpenStudy (anonymous):

do you know how to calculate the length of a vector?

OpenStudy (anonymous):

ermmm. I want to say sqrt(a^2+b^2) but I don't think that's right

OpenStudy (anonymous):

That is correct.

OpenStudy (anonymous):

The best way to remember it is \(\sqrt{v\cdot v}\).

OpenStudy (anonymous):

So in our case \[ \sqrt{a^2+b^2} \]

OpenStudy (anonymous):

Now we are saying: \[ 1 =\sqrt{a^2+b^2} \]

OpenStudy (anonymous):

We square both sides: \[ 1^2 = a^2+b^2 \]

OpenStudy (anonymous):

We already determined that \(a=0\) so \(a^2=0\).

OpenStudy (anonymous):

\[ 1 = b^2 \]

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

so one of the perpendicular vectors is <0,1>?

OpenStudy (anonymous):

Yes, that is true... But you haven't solved for \(b\)! What did you get?

OpenStudy (anonymous):

It is quadratic. It could have 2 solutions.

OpenStudy (anonymous):

I am so confused...

OpenStudy (anonymous):

Okay... let me put it this way...\[ 1=x^2 \]So\[ 0 = x^2-1 \]

OpenStudy (anonymous):

okay, that makes more sense.

OpenStudy (anonymous):

So as you can see, there are two values for \(b\)! That is how come we get two possible unit vectors.